# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ Abstract
However, these previous studies have predominately favored a classical description of the lattice degrees of freedom, which break down for the $5d^{1}$ DPs. For instance, a resonant inelastic x-ray scattering (RIXS) study of $5d^{1}$ DPs at the $L_{3}$ edge finds an asymmetric $j_{\text{eff}}=1/2$ peak [13], which is characteristic of vibrational transitions between different electronic states described by the Franck-Condon principle. Crucially, the asymmetric shape of the $j_{\text{eff}}=1/2$ peak remains unchanged by structural transitions, thereby ruling out the conventional classical mechanisms as an explanation.
The missing piece in the former studies is the quantum nature of lattice vibrations which play a decisive role in shaping the low-energy spectra. In particular, the strong spin-orbit coupling in $5d$ ions enhances the quantum entanglement between the $j_{\text{eff}}=3/2$ states on metal sites, with the quantized vibrations of the surrounding lattice (dynamic JT effect [14]). The experimental fingerprints of these orbital-lattice entangled vibronic states are additional transitions, which emerge due to the dressing of the ground ($j_{\mathrm{eff}}$=3/2) and excited ($j_{\mathrm{eff}}$=1/2) manifold by lattice vibrations that cannot simply be explained by adopting a purely classical description.
@@
where $\lambda$ is the spin-orbit coupling parameter, and $\hat{\bm{l}}$ and $\hat{\bm{s}}$ are the effective orbital and spin angular momenta, respectively. Correspondingly, the energy eigenstates of $\hat{H}_{\text{SO}}$ are the $j_{\text{eff}}=1/2$ $(\Gamma_{7})$ and $j_{\text{eff}}=3/2$ $(\Gamma_{8})$ multiplet states with the eigenvalues of $\lambda$ and $-\lambda/2$, respectively.
@@
via $g$, the dimensionless vibronic coupling parameter between the $d$-$p$ hybridized electronic orbitals $|\gamma\rangle$ ($\gamma=yz,zx,xy$) and the lattice distortion modes $\hat{q}_{\gamma}$ ($\gamma$=$x^{2}$, $y^{2}$, $z^{2}$) with vibration frequency of $\omega$. Here, $\hat{p}_{\gamma}$ is the dimensionless conjugate momenta. Throughout the work, we consider only the $E_{g}$ vibrational modes [Fig. 1(a)] among the vibronically active modes because their coupling strength is the most dominant in $5d^{1}$ DPs [15, 13].
@@
Figure 2: The calculated RIXS spectrum at the (a) rhenium $L_{3}$ and (b) oxygen $K$ edge based on a classical ($g$=0) and quantum ($g$=1.2) description of the lattice degrees of freedom are denoted by the dashed blue and solid red lines, respectively. The faint lines correspond to calculations for intermediate coupling strengths $g$=0.3, 0.6 and 0.9.
@@
To illustrate the spectroscopic signatures of these vibronic transitions, we plot in Fig. 2 the calculated RIXS spectra for a range of coupling strengths ($g$), for an undistorted Ba2CaReO6. We employed the Kramers-Heisenberg formula [27, 28] and the dipole approximation to calculate the spectra that arise from transitions between the vibronic states (see computational details). We applied the fast collision approximation [29, 30] for Re $L_{3}$-edge spectra due to the short lifetime in the intermediate states [31]. We explicitly considered the intermediate $d^{2}$ states with Hund coupling of 0.26 eV [32] for O $K$-edge spectra.
@@
In the classical regime ($g$=0), where the orbital and lattice degrees of freedom are decoupled, our calculations reveal a single, purely-electronic $j_{\mathrm{eff}}$=1/2 transition at $\Delta E$$\sim$0.5 eV. This transition is characterized by a symmetric peak at both the rhenium $L_{3}$ and oxygen $K$ edges. In contrast, when a quantum description of the lattice is adopted ($g>0$), the orbital and lattice degrees of freedom become entangled. For instance, with $g$=1.2, the $j_{\mathrm{eff}}$=1/2 transition broadens asymmetrically and acquires a long tail that extends towards high energies at both the Re $L_{3}$ and O $K$ edges. This broadening is due to the presence of many additional quantum vibronic transitions, which come from the dressing of the $j_{\mathrm{eff}}$=3/2 manifold by lattice vibrations.
@@
Furthermore, a strong asymmetric peak also appears at $\Delta E$$\sim$ 0.15 eV of the O $K$ RIXS spectra for $g$=1.2 [Fig. 2(b)]. The peak originates from a multitude of quantum vibronic transitions between 0-0.4 eV, which appear due to the dressing of the $j_{\mathrm{eff}}$=3/2 ground states by the lattice degrees of freedom and the difference in the dipole transition paths compared with the $L_{3}$-edge RIXS. As shown in Fig. 2(b), such a peak is absent in the classical case, where $g$=0. Therefore, the presence of such a peak is the key spectroscopic fingerprint of quantum vibronic states arising from spin-orbit-lattice entanglement.
@@
Figure 3: Dynamic JT effect in the elastic field. (a)-(c) The APES’s with varied effective elastic field strengths. (d) The evolution of the vibronic energy levels (red dots) with respect to the elastic field arising for tetragonal deformation. The points at $F=0$ correspond to the vibronic excitation energies shown in Fig. 2(b). $F>0$ ($<0$) corresponds to the tetragonal compression (elongation) in eV. The solid blue and dashed green lines correspond field dependence of the harmonic oscillator frequencies, which are obtained from the parabolic fits depicted on the APES.
@@
Next, we consider whether the spin-orbit-lattice entanglement is robust against the elastic field in the case of $5d^{1}$ DP materials. By using realistic interaction parameters for the rhenium-based DPs of $\lambda=0.32$ eV, $g=1.3$, and $\omega=66$ meV [33] (see computational details), we find that the ground vibronic state is stabilized by 30 meV, which is more than two times larger than the static JT stabilization energy of 14 meV and is three times larger than the effective field of the Re compounds measured by the structural transition temperatures. Therefore, given that the size of $F$ is very small ($\sim$ 10 meV), our calculations indicate that the vibronic entanglement effect in cubic $5d^{1}$ DPs should persist strongly, even below the structural transition temperature.
@@
We calculated the Re $L_{3}$-edge RIXS spectra by adopting the interaction parameters described above and also accounting for the experimental configuration shown in the insert of Fig. 4(a). Our calculations reveal that the vibronic coupling modulates the peak shape of the excited $j_{\text{eff}}$=$1/2$ multiplet state at 0.52 eV considerably, resulting in additional transitions at 0.59, 0.67, 0.73 eV. The calculated transition energies and intensities between the vibronic states, denoted by the grey bars in Fig. 4(a), are in good agreement with the experimental data. This nice correspondence between the two is remarkable given that the calculations were based on an undistorted ReO6 octahedron (i.e. for Ba2CaReO6 above $T_{\mathrm{s}}$), in contrast to the measurements which were performed at $T$ = 7 K. Hence, for a more representative calculation of the transitions, we have to account for the effect of the structural distortions on the vibronic level, as depicted in Fig. 3.
@@
Similarly, we calculated the oxygen $K$-edge RIXS spectra, taking into account the experimental configuration shown in the insert of Fig. 4(b). The grey bars indicate the vibronic transitions corresponding to an undistorted ReO6 octahedron at 200 K. Our calculations reveal peaks at 0.52, 0.59, 0.67, 0.73 eV, which arise due to the dressing of the excited states in the $j_{\text{eff}}=1/2$ manifold due to the vibronic coupling.
Our calculation also shows many vibronic transitions in the energy range of 0.05-0.4 eV, arising from the dressing of the phonon degrees of freedom by the electronic $j_{\text{eff}}=3/2$ ground state manifold. As such, these transitions explain that the presence of the inelastic peak at 0.1 eV in the experimental spectrum [Fig. 4(b)], along with its long tail that extends to high energies, originates from vibronic states. Crucially, the experimental and calculated RIXS spectrum at the oxygen $K$-edge represents direct evidence for the presence of orbital-lattice entanglement in Ba2CaReO6.
Next, when the cooperative structural deformation is accounted for, we also find that the distribution of the intensities of the O $K$-edge RIXS spectra does not change considerably. The black lines in Fig. 4(b), denote the calculated transitions at $T$=17 K, based on the distorted structure described earlier. Our experimental data and calculations both demonstrate clearly that the vibronic states are impervious to the quadrupolar ordering in Ba2CaReO6. Therefore, the orbital-lattice entanglement and, hence, the dynamic JT effect persist.
@@ Appendix A Computational details
To include the mean-field effect into the vibronic states, we numerically diagonalized the above Hamiltonian with Eq. (6). We determined $F_{z^{2}}=3.3$ meV so that the expectation value of $\hat{q}_{z^{2}}$ corresponds to the experimental structure at 17 K [21].
We consider the RIXS processes involving the transitions between the Re $2p$/O $1s$ to the $5d$ $t_{2g}$ orbitals. For the calculations of the RIXS intensities, we employed the Kramers-Heisenberg formula [27, 28].
@@
×$\times\delta\left(E_{\nu^{\prime}}+\hslash\omega_{k^{\prime}}-E_{\nu}-\hslash\omega_{k}\right),$ (7)
where $\hat{P}_{\bm{k}\alpha}$ is the dipole moments along the polarization vector of the light, and $\rho_{\nu}$ is the canonical distribution for single $5d^{1}$ site. The initial and final states denoted by $\nu,\nu^{\prime}$ in Eq. (7) are the vibronic types, and the intermediate states $\mu$ are $5d^{2}$ multiple type. The broadening of the intermediate states $\Gamma_{\mu}$ is 150 meV [18]. In the Re $L_{3}$-edge RIXS calculations, we ignored the multiplet structure of the intermediate states as in Ref. [13]. In the O $K$-edge RIXS, the transition occurs from the O $1s$ orbital to the Re 5d-like $t_{2g}$ orbitals hybridized with the O $2p$ orbitals. To reproduce the RIXS spectra, we convoluted the intensities (7) with the Lorentzian function: The linewidths are $\Gamma=115$ meV for Re $L_{3}$- and $\Gamma=35$ meV for O $K$-edge RIXS spectra, respectively. We calculated only the inelastic contributions for the O $K$-edge RIXS spectra.

### v3

@@ Abstract
However, these previous studies have predominately favored a classical description of the lattice degrees of freedom, which break down for the $5d^{1}$ DPs. For instance, a resonant inelastic x-ray scattering (RIXS) study of $5d^{1}$ DPs at the $L_{3}$ edge finds an asymmetric $j_{\text{eff}}=1/2$ peak [13]. Crucially, the asymmetric shape of the $j_{\text{eff}}=1/2$ peak remains unchanged by structural transitions, thereby ruling out the conventional Franck-Condon mechanism as an explanation.
The missing piece in many of the former studies is the quantum nature of lattice vibrations which play a decisive role in shaping the low-energy spectra. In particular, the strong spin-orbit coupling in $5d$ ions enhances the quantum entanglement between the $j_{\text{eff}}=3/2$ ground state on metal sites, with the quantized vibrations of the surrounding lattice (dynamic JT effect [14]). The experimental fingerprints of these orbital-lattice entangled vibronic states are additional transitions, which emerge due to the dressing of the $j_{\mathrm{eff}}$=3/2 manifold by lattice vibrations that cannot simply be explained by adopting a purely classical description. Indeed, the transitions between the vibronic ground states and excited $j_{\text{eff}}=1/2$ states make the $j_{\text{eff}}=1/2$ RIXS peak of the $5d^{1}$ DPs asymmetric [13].
@@
where $\lambda$ is the spin-orbit coupling parameter, and $\hat{\bm{l}}$ and $\hat{\bm{s}}$ are the effective orbital and spin angular momenta, respectively. Correspondingly, the energy eigenstates of $\hat{H}_{\text{SO}}$ are the $j_{\text{eff}}=1/2$ $(\Gamma_{7})$ and $j_{\text{eff}}=3/2$ $(\Gamma_{8})$ multiplet states with the eigenvalues of $\lambda$ and $\lambda/2$, respectively.
@@
via $g$, the dimensionless vibronic coupling parameter between the $d$-$p$ hybridized electronic orbitals $|\gamma\rangle$ ($\gamma=yz,zx,xy$) and the lattice distortion modes $\hat{q}_{\gamma}$ ($\gamma$$x^{2}$, $y^{2}$, $z^{2}$) with vibration frequency of $\omega$. Here, $\hat{p}_{\gamma}$ is the dimensionless conjugate momenta. Throughout the work, we consider only the $E_{g}$ vibrational modes [Fig. 1(a)] among the vibronically active modes because their coupling strength is the most dominant in $5d^{1}$ DPs [15, 13].
@@
Figure 2: The calculated RIXS spectrum at the (a) rhenium $L_{3}$ and (b) oxygen $K$ edge based on a classical (gg=0) and quantum (gg=1.3) description of the lattice degrees of freedom are denoted by the dashed blue and solid red lines, respectively. The faint lines correspond to calculations for intermediate coupling strengths gg=0.325, 0.650 and 0.975.
@@
To illustrate the spectroscopic signatures of these vibronic transitions, we plot in Fig. 2 the calculated RIXS spectra for a range of coupling strengths ($g$), for an undistorted Ba2CaReO6. We employed the Kramers-Heisenberg formula [27, 28] and the dipole approximation to calculate the spectra that arise from transitions between the vibronic states (see computational details).
@@
In the classical regime (gg=0), where the orbital and lattice degrees of freedom are decoupled, our calculations reveal a single, purely-electronic $j_{\mathrm{eff}}$=1/2 transition at $\Delta E\sim$0.5 eV. This transition is characterized by a symmetric peak at both the rhenium $L_{3}$ and oxygen $K$ edges. In contrast, when a quantum description of the lattice is adopted ($g>0$), the orbital and lattice degrees of freedom become entangled. For instance, with gg=1.3, the $j_{\mathrm{eff}}$=1/2 transition broadens asymmetrically and acquires a long tail that extends towards high energies at both the Re $L_{3}$ and O $K$ edges. This broadening is due to the presence of many additional quantum vibronic transitions, which come from the dressing of the $j_{\mathrm{eff}}$=3/2 manifold by lattice vibrations.
@@
Furthermore, a strong asymmetric peak also appears at $\Delta E\sim$ 0.15 eV of the O $K$ RIXS spectra for gg=1.3 [Fig. 2(b)]. The peak originates from a multitude of quantum vibronic transitions between 0-0.4 eV. As shown in Fig. 2(b), such a peak is absent in the classical case, where gg=0. Therefore, the presence of such a peak is the key spectroscopic fingerprint of quantum vibronic states arising from spin-orbit-lattice entanglement.
Vibronic transitions are more pronounced in the RIXS spectra at the O $K$ edge than at the Re $L_{3}$ edge. This arises because the $K$-edge scattering amplitude contains both magnetic and quadrupolar contributions [29], whereas the $L_{3}$-edge amplitude is purely quadrupolar—without spin-flip processes when intermediate-energy structures are neglected. Moreover, the larger quadrupolar component at the $K$ edge enhances vibronic excitations, as orbital–lattice entanglement is also time-even in nature.
@@
Figure 3: Dynamic JT effect in the elastic field. (a)-(c) The APES’s with varied effective elastic field strengths. (d) The evolution of the vibronic energy levels (red dots) with respect to the elastic field arising for tetragonal deformation. The points at $F=0$ correspond to the vibronic excitation energies shown in Fig. 2(b). $F>0$ (<0<0) corresponds to the tetragonal compression (elongation) in eV. The solid blue and dashed green lines correspond field dependence of the harmonic oscillator frequencies, which are obtained from the parabolic fits depicted on the APES.
@@
Next, we consider whether the spin-orbit-lattice entanglement is robust against the elastic field in the case of $5d^{1}$ DP materials. By using realistic interaction parameters for the rhenium-based DPs of $\lambda=0.312$ eV, $g=1.3$, and $\omega=66$ meV [30] (see computational details), we find that the ground vibronic state is stabilized by 30 meV, which is more than two times larger than the static JT stabilization energy of 12 meV and is three times larger than the effective field of the Re compounds measured by the structural transition temperatures. Therefore, given that the size of $F$ is very small ($\sim$ 10 meV), our calculations indicate that the vibronic entanglement effect in cubic $5d^{1}$ DPs should persist strongly, even below the structural transition temperature.
@@
We calculated the Re $L_{3}$-edge RIXS spectra by adopting the interaction parameters described above and also accounting for the experimental configuration shown in the insert of Fig. 4(a). Our calculations reveal that the vibronic coupling modulates the peak shape of the excited $j_{\text{eff}}$=$1/2$ multiplet state at 0.52 eV considerably, resulting in additional transitions at 0.59, 0.67, 0.73 eV. The calculated transition energies and intensities between the vibronic states, denoted by the grey bars in Fig. 4(a), are in good agreement with the experimental data. This correspondence between the two is remarkable given that the calculations were based on an undistorted ReO6 octahedron (i.e. for Ba2CaReO6 above $T_{\mathrm{s}}$), in contrast to the measurements which were performed at $T$ = 7 K. Hence, for a more representative calculation of the transitions, we have to account for the effect of the structural distortions on the vibronic level, as depicted in Fig. 3.
@@
Similarly, we calculated the oxygen $K$-edge RIXS spectra with the same model as for the $L_{3}$-edge RIXS, taking into account the experimental configuration shown in the insert of Fig. 4(b). The grey bars indicate the vibronic transitions corresponding to an undistorted ReO6 octahedron at 200 K. Our calculations reveal peaks at 0.52, 0.59, 0.67, 0.73 eV, which arise due to the vibronic transitions to the $j_{\text{eff}}=1/2$ manifold with vibrational excitations.
Our calculation also shows many transitions between the vibronic states from the $j_{\text{eff}}=3/2$ manifold in the energy range of 0.05-0.4 eV. As such, these transitions explain that the presence of the inelastic peak at 0.1 eV in the experimental spectrum [Fig. 4(b)], along with its long tail that extends to high energies, originates from vibronic states: The peak originates from the transitions to the vibronic quartet states at 0.067 eV and 0.104 eV. Crucially, the experimental and calculated RIXS spectrum at the oxygen $K$-edge represents direct evidence for the presence of orbital-lattice entanglement in Ba2CaReO6.
Next, when the cooperative structural deformation is accounted for, we also find that the distribution of the intensities of the O $K$-edge RIXS spectra does not change considerably. The black lines in Fig. 4(b), denote the calculated transitions at TT=17 K, based on the distorted structure described earlier. Our experimental data and calculations both demonstrate clearly that the vibronic states are impervious to the quadrupolar ordering in Ba2CaReO6. Therefore, the orbital-lattice entanglement and, hence, the dynamic JT effect persist.
@@ Appendix A Computational details
To include the mean-field effect on the vibronic states, we numerically diagonalized the above Hamiltonian with Eq. (6). We determined $F_{z^{2}}=3.3$ meV so that the expectation value of $\hat{q}_{z^{2}}$ corresponds to the experimental structure at 17 K [21].
We consider the RIXS processes involving the transitions between the Re 2p2p/O $1s$ to the $5d$ $t_{2g}$ orbitals. For the calculations of the RIXS intensities, we employed the Kramers-Heisenberg formula [27, 28].
@@
$\times\delta\left(E_{\nu^{\prime}}+\hslash\omega_{k^{\prime}}-E_{\nu}-\hslash\omega_{k}\right),$ (7)
where $\hat{P}_{\bm{k}\alpha}$ is the dipole moments along the polarization vector of the light, and $\rho_{\nu}$ is the canonical distribution for single $5d^{1}$ site. The initial and final states denoted by $\nu,\nu^{\prime}$ in Eq. (7) are the vibronic types, and the intermediate states $\mu$ are $5d^{2}$ vibronic states with a core-hole. We used the Hund coupling of 0.26 eV [33]. The broadening of the intermediate states $\Gamma_{\mu}$ is 2.25 eV for Re $L_{3}$-edge [34] and 150 meV for O $K$-edge RIXS [18]. In the O $K$-edge RIXS, the transition occurs between the $p$-type molecular orbitals, each of which is composed of two O $1s$ orbitals, and the $t_{2g}$ orbitals. The $K$ edge scattering amplitude is the same as the sum of the $L_{2}$ and $L_{3}$ ones with the degenerate core-hole states. To reproduce the RIXS spectra, we convoluted the intensities (7) with the Lorentzian function: The full widths at half maxima are 0.13 eV for Re $L_{3}$- and 0.1 eV for O $K$-edge RIXS spectra. We calculated only the inelastic contributions for the O $K$-edge RIXS spectra. We scaled the transition intensities within the $j_{\text{eff}}=3/2$ manifold by 0.445 times to account for the energy dependent absorption of low-energy photon. In the ordered phases, we averaged the spectra over the domains of the quadrupolar phases.